Trucks in a delivery fleet travel a mean of 80 miles per day with a standard deviation of 30 miles per day. The mileage per day is distributed normally. Find the probability that a truck drives between 97 and 107 miles in a day. (Round your answer to 4 decimal places)
Answer #1

Solution :

Given that ,

mean = = 80

standard deviation = =30

P(97< x < 107) = P[(97-80) /30 < (x - ) / < (107-80) / 30)]

= P(0.57 < Z <0.9 )

= P(Z <0.9 ) - P(Z < 0.57)

Using z table   

= 0.8159-0.7157

probability= 0.1002